OpenStudy (anonymous):

0.57142857142 as a fraction

OpenStudy (ageta):

571428571/1000000000

OpenStudy (ageta):

Work: Write the decimal number as a fraction (over 1) 0.571428571 = 0.571428571 / 1 Multiplying by 1 to eliminate 9 decimal places we multiply top and bottom by 9 10's Numerator (N) N = 0.571428571 × 10 × 10 × 10 × 10 × 10 × 10 × 10 × 10 × 10 = 571428571 Denominator (D) D = 1 × 10 × 10 × 10 × 10 × 10 × 10 × 10 × 10 × 10 = 1000000000 N / D = 571428571 / 1000000000 Simplifying our fraction = 571428571/1000000000 = 571428571/1000000000

OpenStudy (ageta):

if you have a lot of question go to http://www.calculatorsoup.com/calculators/math/decimal-to-fraction-calculator.php

OpenStudy (anonymous):

i bet it is something repeating, maybe \(\frac{1}{7}\)

OpenStudy (anonymous):

that was a bad guess actually \(\frac{4}{7}\) http://www.wolframalpha.com/input/?i=4%2F7

OpenStudy (ageta):

if you simplify it Why divide by -2.14748e+09? -2.14748e+09 is the Greatest Common Divisor (GCD) or Greatest Common Factor (GCF) of the numbers 5.71429e+10 and 1e+11. So, this fraction reduced to lowest terms is So your final answer is: 0.57142857142 can be written as the fraction

OpenStudy (anonymous):

i still think it is \(\frac{4}{7}\)

OpenStudy (ageta):

4/7 ya he is right